Reginald is expecting steady growth of 10 percent per year in profits from his new company. All profits are going to be invested at 20 percent interest. If profits for this year (at the end of the year) total 10000$, how much will be saved at the end of 10 years ?

Answer : 35981.09 $

Can you please try to find the Present Value first, then the Future Value ? Thank you
Answer #1

First find the present value using geometric gradient where A = 10000, n = 10, g = 10%, i = 20%

P A1-(1+g)(1+ i) -n 8-1

PV = 10,000 x (1 - (1.10^10)*(1.20^-10))/(0.20 - 0.10)

= $58,109.61

Now find the future value of this money which is 58,109.61*(1 + 20%)^10 = $359,799.40 or approximately $359,800.

Using growth adjusted rate of interest, we have i = (1 + 20%)/(1 + 10%) - 1 = 9.09%

Present value PV = 10,000 x (P/A, 9.09%, 10) / (1 + 10%) = 10,000 x 6.3923 / 1.1 = 58,111.82

Future value = 58,111.82(F/P, 20%, 10)

= 58,111.82*6.1917

= 359,810.9
Answer #2

fvuck you nigger

source: asfsa
answered by: anonymous